Peter decides to mix grades of gasoline in his truck. He put in 6 gallons of regular in 7 gallons of premium or total cost of $34.09. If premium gasoline cost $.19 more per gallon then regular what was the price of each grade of gasoline
Answer by Boreal(15235)

6R+7P=34.09
P=R+0.19
substitute
6R+7(R+0.19)=34.09
6R+7R+1.33=34.09
13R=32.76
R=$2.52 *6=$15.12
P=$2.71 *7=$18.97
They add to $34.09